OpenStudy (anonymous):

2sinx -2cosx = √2, solve for x. ANSWER: 5π/12 and 13π/12

hartnn (hartnn):

divide by sqrt 2 on both sides what u get?

OpenStudy (anonymous):

Haha, that'll give me 2/√2(sinx) - 2/√2(cosx) = 1

OpenStudy (anonymous):

Then I'd factor √2/2 out and eventually end up with the same wrong answer I've been getting...

OpenStudy (anonymous):

sorry, 2/√2, not √2/2

hartnn (hartnn):

note that 1/sqrt 2 = sin 45 = cos 45 so u have cos 45 sin x - sin 45 cos x you remember sin(A_B) formula ?

OpenStudy (anonymous):

Haha, it isn't 1/√2 though, it's 2/√2?

OpenStudy (anonymous):

Ah, but I guess that's the same as 2(1/√2)...so then is it 2(sin45-cos45)=1?

hartnn (hartnn):

yeah, youu can factor that 2 out

hartnn (hartnn):

where did 'x' go ? :O

OpenStudy (anonymous):

Haha, yeah that's what I was wondering.

OpenStudy (anonymous):

xD

hartnn (hartnn):

sin A cos B - cos A sin B = sin (A-B) cos 45 sin x - sin 45 cos x = sin (x-45) makes sense ?

OpenStudy (anonymous):

Not at all.

hartnn (hartnn):

did u know this ? sin A cos B - cos A sin B = sin (A-B)

OpenStudy (anonymous):

Why are you introducing sin and cos of 45 degrees? Yeah, I know the double angle identities.

OpenStudy (anonymous):

Or compound, whichever one this is.

hartnn (hartnn):

to simplify the left side 1/sqrt 2 = sin 45 = cos 45 that simplifies left side greatly

OpenStudy (anonymous):

This doesn't make sense. Thanks for trying, but it's only confusing me further.

hartnn (hartnn):

so did u not get this ? sin A cos B - cos A sin B = sin (A-B) sin x cos 45 - cos x sin 45 = sin (x-45)

myininaya (myininaya):

@KinzaN , @hartnn has a good approach

myininaya (myininaya):

Is there anything that he said you can point out that doesn't make sense so I can clarify it or so he can?

myininaya (myininaya):

\[2 ( \sin(x)-\cos(x))=\sqrt{2} \] \[\frac{1}{\sqrt{2}} \sin(x)-\frac{1}{\sqrt{2}} \cos(x)=\frac{1}{2}\] Have you gotten this far?

OpenStudy (anonymous):

Yeah, but beyond that I get stuck after getting an answer of 5π/12-- and I also don't understand why this is the only method that works. There must be another that's simpler which we're expected to use.

OpenStudy (anonymous):

Sine is positive in the first and second quadrants, so why is their answer of 13π/12 in the third? Wah.

OpenStudy (anonymous):

Ah, never mind! I got both solutions using this method. TYTY.